# Robust Unsupervised Crowd Counting and Localization with Adaptive Resolution SAM

### v1

@@ Abstract
The existing crowd counting models require extensive training data, which is time-consuming to annotate. To tackle this issue, we propose a simple yet effective crowd counting method by utilizing the Segment-Everything-Everywhere Model (SEEM), an adaptation of the Segmentation Anything Model (SAM), to generate pseudo-labels for training crowd counting models. However, our initial investigation reveals that SEEM’s performance in dense crowd scenes is limited, primarily due to the omission of many persons in high-density areas. To overcome this limitation, we propose an adaptive resolution SEEM to handle the scale variations, occlusions, and overlapping of people within crowd scenes. Alongside this, we introduce a robust localization method, based on Gaussian Mixture Models, for predicting the head positions in the predicted people masks. Given the mask and point pseudo-labels, we propose a robust loss function, which is designed to exclude uncertain regions based on SEEM’s predictions, thereby enhancing the training process of the counting network. Finally, we propose an iterative method for generating pseudo-labels. This method aims at improving the quality of the segmentation masks by identifying more tiny persons in high-density regions, which are often missed in the first pseudo-labeling stage. Overall, our proposed method achieves the best unsupervised performance in crowd counting, while also being comparable to some classic supervised fully methods. This makes it a highly effective and versatile tool for crowd counting, especially in situations where labeled data is not available.
@@ 1 Introduction
To tackle this challenge, we introduce a robust unsupervised method that utilizes the Segmentation Anything Model (SAM) [16] to generate pseudo labels. However, SAM is not able to predict semantic labels. Therefore, the Segment-Everything-Everywhere Model (SEEM) [48] is utilized to predict person masks. Our findings indicate that using SEEM directly is not effective, since it often misses people due to occlusions and overlapping (see 1024x1024 image in Fig. 1), which is due to the limited availability of dense crowd images in its training data. To address this, we propose an adaptive resolution SEEM (AdaSEEM) that can zoom in on areas of high density as needed. This enhancement allows for more precise segmentation of smaller persons in crowded regions as shown in Figure 1. In addition, we propose a robust head localization method to estimate the head locations accurately by modeling the mask distribution as a Gaussian Mixture Model (GMM), enabling the generation of more effective point pseudo-labels.
@@
Finally, to enhance performance, we adopt an iterative approach for generating pseudo masks, using the point predictions from the well-trained counting network as prompts for AdaSEEM. This helps in identifying missing individuals in high-density areas. Once these new masks are created, they are fused with those from the previous stage to create a more comprehensive and accurate set of pseudo-labels. Subsequently, we employ the same methodology to estimate head point pseudo-labels within these updated masks. With these refined masks and head locations, we proceed to train the counting networks, thereby improving their accuracy and reliability in densely populated scenes.
@@ 2.1 Supervised Methods
Recent crowd counting research has predominantly focused on deep learning, with significant improvements achieved through training with extensive labeled data [47]. Innovations in network structures [1, 8] and the development of various loss functions [40] have enhanced performance and robustness. Kang and Chan [15] introduced the use of an image pyramid to address scale variation. Further enhancements include the exploitation of contextual information [31, 44] and the development of cross-scene crowd counting methods to improve generalization [46]. Wang et al. [41] proposed the use of a synthetic dataset, while others have explored the use of correlation information to boost generalization capabilities [36, 43]. Innovative approaches in loss function design, such as learnable density maps for enhanced supervision, have been proposed [34, 37]. Direct use of point annotations during training has shown improved counting and localization [40, 33, 26, 23], and robust loss functions have been developed to address annotation noise [35, 39]. Recently, Transformer-based methods have demonstrated exceptional performance in both crowd counting and localization [21, 19].
@@ 2.2 Semi-Supervised And Unsupervised Methods
To alleviate the burden of extensive annotation, several innovative approaches have been proposed in the realm of crowd counting [42]. Change Loy et al. [6] suggest the use of unlabeled videos, thereby reducing the dependency on fully labeled datasets. Meng et al. [27] introduce a method to model spatial uncertainty, enhancing the efficacy of semi-supervised counting. The concept of training models with partial annotations has also been explored [45], offering a practical alternative to fully supervised methods. Furthermore, a supervised uncertainty estimation strategy is presented in [17], providing a novel approach to address annotation challenges. Additionally, the use of optimal transport minimization [22] has been proposed for crowd localization in semi-supervised settings, further contributing to the development of more efficient and less labor-intensive methods in the field of crowd counting.
@@ 3.2 Robust Localization for Point Pseudo-labels
3.3 Counter Training
@@
Method Year Label UCF-QNRF JHU ShTech A ShTech B
MAE MSE MAE MSE MAE MSE MAE MSE
Zhang et al. [46] CVPR 15 Point - - - - 181.8 277.7 32.0 49.8
MCNN [47] CVPR 16 Point 277.0 426.0 188.9 483.4 110.2 173.2 26.4 41.3
Switch CNN [1] CVPR 17 Point 228.0 445.0 - - 90.4 135.0 21.6 33.4
LSC-CNN [29] TPAMI 21 Point 120.5 218.2 112.7 454.4 66.4 117.0 8.1 12.7
SDA+DM [25] ICCV 21 Point 80.7 146.3 59.3 248.9 55.0 92.7 - -
CLTR [19] ECCV 22 Point 85.8 141.3 59.5 240.6 56.9 95.2 6.5 10.2
MAN [21] CVPR 22 Point 77.3 131.5 53.4 209.9 56.8 90.3 - -
Chfl [30] CVPR 23 Point 80.3 137.6 57.0 235.7 57.5 94.3 6.9 11.0
STEERER [8] ICCV 23 Point 74.3 128.3 54.3 238.3 54.5 86.9 5.8 8.5
SFCN [41] CVPR 19 Point (GCC [41]) 275.5 458.5 - - 160.0 216.5 22.8 30.6
RCC [10] arXiv 22 Point (FSC [28]) - - - - 240.1 366.9 66.6 104.8
CLIP-Count [14] arXiv 23 Point (FSC [28]) - - - - 192.6 308.4 45.7 77.4
CSS-CNN-Random [2] ECCV 22 None 718.7 1036.3 320.3 793.5 431.1 559.0 - -
Random* - None 633.6 978.9 297.5 801.6 411.1 511.1 158.7 287.4
CSS-CNN [2] ECCV 22 None 437.0 722.3 217.6 651.3 179.3 295.9 - -
CrowdCLIP [20] CVPR 23 None 283.3 488.7 213.7 576.1 146.1 236.3 69.3 85.8
Ours (Stage 1) None 195.9 343.0 109.5 428.7 125.4 226.7 34.4 55.2
Ours (Stage 2) None 181.2 304.7 105.1 390.5 122.8 217.8 33.3 53.2
Ours (Stage 3) None 182.3 289.9 102.7 360.7 102.6 176.3 35.6 51.7
@@ 3.4 Iterative Pseudo-label Generation
     # Train Stage-$k$ counter
@@
Figure 3: The masks generated from different methods. From left to right are: SEEM, adaptive resolution SEEM (AdaSEEM), and AdaSEEM + Stage 1 predictions. In (c), the new pseudo-label masks that were found in Stage 1 are highlighted with blue ellipses.
@@ 4.1 Experimental Settings
Dataset: We evaluate the proposed method on JHU-CROWD dataset [32], UCF-QNRF [13], and ShanghaiTech [47] datasets. The JHU-CROWD dataset is a comprehensive large-scale dataset, comprising 4,371 images. It is divided into three subsets: 2,722 images for training, 500 for validation, and 1,600 for testing. The UCF-QNRF dataset includes 1,535 images, with 1,201 designated for training and 334 for testing. The ShanghaiTech dataset is split into two parts: the ShanghaiTech A, which contains a total of 782 images, divided into 482 for training and 300 for testing, and the ShanghaiTech B, which includes 1,116 images, with 716 used for training and 400 for testing.
@@
The parameters $\omega$ and $\beta$ in our loss function play crucial roles in optimizing performance. These parameters are set to 100 and 0.01, respectively, based on the ablation study in Figures 10 and 10. The threshold $\tau$ in AdaSEEM is set to 0.3 according to the experimental result shown in Figure 10.
@@ 4.2 Comparison with State-of-the-art Methods
To assess the effectiveness of our proposed method, we conducted a thorough evaluation by comparing it with both state-of-the-art unsupervised and supervised methods. The results of this comparison are detailed in Table 1. First, our proposed method outperforms other unsupervised methods in terms of MAE and MSE, and the margin of improvement is significant. This underscores the effectiveness of our approach in addressing the challenges inherent to unsupervised counting. Second, the comparison also reveals that the performance in Stage 2 of our method is better than in Stage 1 across all datasets. This improvement validates the effectiveness of our iterative pseudo-label generation strategy. By refining the pseudo-labels, the model is able to achieve more accurate and reliable counting results. We also compare the proposed method with cross-domain methods, which train on a source dataset and test on the target dataset, in Table 1 and achieve superior performance for most of the cases. Finally, the proposed method also compares favorably with some classic supervised methods. However, there is still considerable room for improvement, particularly in handling densely crowded datasets.
@@ 4.3 Ablation Study
Adaptive resolution SAM We conducted counting experiments to evaluate the effectiveness of AdaSEEM, and the results are presented in Table 2. The performance of SEEM on its own was found to be the least effective, indicating that directly using SEEM is not optimal due to the omission of many small individuals in high-density areas. However, with the implementation of the proposed adaptive resolution strategy, there was a noticeable improvement in performance, especially for the high-density dataset UCF-QNRF. In Figure 4, we can also observe significant improvement in high-density images when using AdaSEEM. This improvement underscores the efficacy of the adaptive resolution strategy in accurately segmenting small persons in densely populated regions. The effectiveness of this approach is further confirmed in Figure 3(a, b), where more individuals are segmented with the AdaSEEM compared to the base model. These findings collectively highlight the significance of the adaptive resolution strategy in enhancing the segmentation capabilities of SEEM in complex crowd counting scenarios.
@@
Ours Stage 1 training 125.4 226.7 195.9 343.0
AdaSEEM + Stage 1 323.4 470.8 347.5 612.1
Ours Stage 2 training 122.8 217.8 181.2 304.7
@@
Figure 9: Error v.s. stages.
Figure 6: Error v.s. $\omega$.
Figure 7: Error v.s. $\beta$.
Figure 8: Error v.s. $\gamma$.
Figure 9: Error v.s. stages.
@@
Iterative pseudo-label generation To enhance mask quality, we introduce an iterative pseudo-labels generation approach, leveraging the trained counting network. First, we predict individual locations in training images, considering local maxima above a threshold as potential person localizations, following [38]. These predicted locations are then used as prompts for SEEM segmentation, effectively localizing new people in dense areas. As Figure 3 illustrates, this method detects more people, evidenced by the increased mask count. Performance comparisons in Table 2 show marked improvements with iterative pseudo-label generation by comparing “AdaSEEM” and “AdaSEEM + Stage 1”. Moreover, a newly trained counting network with these refined masks outperforms the prior stage, confirming the method’s efficacy (Ours Stage 2 training).
To determine the optimal number of iterations, we experimented on ShanghaiTech A, with results depicted in Figure 10. The findings indicate that peak performance is attained at the third stage, after which the performance converges. Consequently, we opted for three stages in subsequent experiments.
Loss hyperparameters Figures 10 and 10 show the ablation studies for different values of the loss hyperparameters, $\omega$ and $\beta$. Figure 10 shows the ablation study for $\tau$.
Localization performance We further evaluate the localization performance of the proposed method on UCF-QNRF. The performance of our proposed unsupervised method was benchmarked against existing supervised methods, filling a gap as there were no comparable unsupervised crowd localization methods. Despite the lack of manual labeling during training, our method demonstrated commendable precision, which outperforms several supervised counterparts, as shown in Table 3. The recall of our method is lower than supervised approaches, but can be improved significantly using the 2nd stage training, which confirms that more missed people are detected and pseudo-labeled. While the overall localization performance of the proposed method is still limited and falls short of the state-of-the-art supervised methods, the results are promising, particularly considering the absence of manual labels.
@@
Ours (Stage 1) None 0.777 0.101 0.456
Ours (Stage 2) None 0.677 0.263 0.476

### v2

# Robust Zero-Shot Crowd Counting and Localization With Adaptive Resolution SAM

@@ Abstract
The existing crowd counting models require extensive training data, which is time-consuming to annotate. To tackle this issue, we propose a simple yet effective crowd counting method by utilizing the Segment-Everything-Everywhere Model (SEEM), an adaptation of the Segmentation Anything Model (SAM), to generate pseudo-labels for training crowd counting models. However, our initial investigation reveals that SEEM’s performance in dense crowd scenes is limited, primarily due to the omission of many persons in high-density areas. To overcome this limitation, we propose an adaptive resolution SEEM to handle the scale variations, occlusions, and overlapping of people within crowd scenes. Alongside this, we introduce a robust localization method, based on Gaussian Mixture Models, for predicting the head positions in the predicted people masks. Given the mask and point pseudo-labels, we propose a robust loss function, which is designed to exclude uncertain regions based on SEEM’s predictions, thereby enhancing the training process of the counting network. Finally, we propose an iterative method for generating pseudo-labels. This method aims at improving the quality of the segmentation masks by identifying more tiny persons in high-density regions, which are often missed in the first pseudo-labeling iteration. Overall, our proposed method achieves the best unsupervised performance in crowd counting, while also being comparable to some classic supervised fully methods. This makes it a highly effective and versatile tool for crowd counting, especially in situations where labeled data is not available.
Keywords:
Crowd Counting Crowd Localization Segment Anything
@@ 1 Introduction
To tackle this challenge, we introduce a robust unsupervised method that utilizes the Segmentation Anything Model (SAM) [20] to generate pseudo labels. However, SAM is not able to predict semantic labels. Therefore, the Segment-Everything-Everywhere Model (SEEM) [55] is utilized to predict person masks. Large foundation models are shown to be useful for downstream tasks [54]. However, our findings indicate that using SEEM directly is not effective, since it often misses people due to occlusions and overlapping (see 1024x1024 image in Fig. 1), which is due to the limited availability of dense crowd images in its training data. To address this, we propose an adaptive resolution SEEM (AdaSEEM) that can zoom in on areas of high density as needed. This enhancement allows for more precise segmentation of smaller persons in crowded regions as shown in Figure 1. In addition, we propose a robust head localization method to estimate the head locations accurately by modeling the mask distribution as a Gaussian Mixture Model (GMM), enabling the generation of more effective point pseudo-labels.
@@
Finally, to enhance performance, we adopt an iterative approach for generating pseudo masks, using the point predictions from the well-trained counting network as prompts for AdaSEEM. This helps in identifying missing individuals in high-density areas. Once these new masks are created, they are fused with those from the previous iteration to create a more comprehensive and accurate set of pseudo-labels. Subsequently, we employ the same methodology to estimate head point pseudo-labels within these updated masks. With these refined masks and head locations, we proceed to train the counting networks, thereby improving their accuracy and reliability in densely populated scenes.
@@ 2.1 Supervised Methods
Recent crowd counting research has predominantly focused on deep learning, with significant improvements achieved through training with extensive labeled data [53, 15, 52, 7, 45]. Innovations in network structures [2, 11, 9] and the development of various loss functions [44] have enhanced performance and robustness. [19] introduced the use of an image pyramid to address scale variation. Further enhancements include the exploitation of contextual information [35, 49, 8] and the development of cross-scene crowd counting methods to improve generalization [51]. [46] proposed the use of a synthetic dataset, while others have explored the use of correlation information to boost generalization capabilities [41, 48]. Innovative approaches in loss function design, such as learnable density maps for enhanced supervision, have been proposed [38, 42]. Direct use of point annotations during training has shown improved counting and localization [44, 37, 30, 27], and robust loss functions have been developed to address annotation noise [39, 43]. Recently, Transformer-based methods have demonstrated exceptional performance in both crowd counting and localization [25, 24].
@@ 2.2 Semi-Supervised And Unsupervised Methods
To alleviate the burden of extensive annotation, several innovative approaches have been proposed in the realm of crowd counting [47]. [6] suggest the use of unlabeled videos, thereby reducing the dependency on fully labeled datasets. [31] introduce a method to model spatial uncertainty, enhancing the efficacy of semi-supervised counting. The concept of training models with partial annotations has also been explored [50], offering a practical alternative to fully supervised methods. Furthermore, a supervised uncertainty estimation strategy is presented in [21], providing a novel approach to address annotation challenges. Additionally, the use of optimal transport minimization [26] has been proposed for crowd localization in semi-supervised settings, further contributing to the development of more efficient and less labor-intensive methods in the field of crowd counting.
@@ 3.2 Robust Localization for Point Pseudo-labels
3.3 Counter Training with Robust Loss
@@
Method Year Label UCF-QNRF JHU ShTech A ShTech B UCF-CC-50
MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
Zhang et al. [51] CVPR 15 Point - - - - 181.8 277.7 32.0 49.8 467.0 498.5
MCNN [53] CVPR 16 Point 277.0 426.0 188.9 483.4 110.2 173.2 26.4 41.3 377.6 509.1
Switch CNN [2] CVPR 17 Point 228.0 445.0 - - 90.4 135.0 21.6 33.4 318.1 439.2
LSC-CNN [33] TPAMI 21 Point 120.5 218.2 112.7 454.4 66.4 117.0 8.1 12.7 225.6 302.7
SDA+DM [28] ICCV 21 Point 80.7 146.3 59.3 248.9 55.0 92.7 - - - -
CLTR [24] ECCV 22 Point 85.8 141.3 59.5 240.6 56.9 95.2 6.5 10.2 - -
MAN [25] CVPR 22 Point 77.3 131.5 53.4 209.9 56.8 90.3 - - - -
Chfl [34] CVPR 23 Point 80.3 137.6 57.0 235.7 57.5 94.3 6.9 11.0 - -
STEERER [11] ICCV 23 Point 74.3 128.3 54.3 238.3 54.5 86.9 5.8 8.5 - -
SFCN [46] CVPR 19 Point (GCC [46]) 275.5 458.5 - - 160.0 216.5 22.8 30.6 487.2 689.0
RCC [13] arXiv 22 Point (FSC [32]) - - - - 240.1 366.9 66.6 104.8 - -
CLIP-Count [18] arXiv 23 Point (FSC [32]) - - - - 192.6 308.4 45.7 77.4 - -
CSS-CNN-Rnd. [1] ECCV 22 None 718.7 1036.3 320.3 793.5 431.1 559.0 - - 1279.3 1567.9
Random* - None 633.6 978.9 297.5 801.6 411.1 511.1 158.7 287.4 1251.6 1497.8
CSS-CNN [1] ECCV 22 None 437.0 722.3 217.6 651.3 179.3 295.9 - - 564.9 959.4
CrowdCLIP [23] CVPR 23 None 283.3 488.7 213.7 576.1 146.1 236.3 69.3 85.8 438.3 604.7
Ours (Iter. 0) None 195.9 343.0 109.5 428.7 125.4 226.7 34.4 55.2 424.5 597.1
Ours (Iter. 1) None 181.2 304.7 105.1 390.5 122.8 217.8 33.3 53.2 382.7 444.5
Ours (Iter. 2) None 182.3 289.9 102.7 360.7 102.6 176.3 35.6 51.7 376.6 578.2
@@ 3.4 Iterative Pseudo-label Generation
     # Train Iteration-$k$ counter
@@
Figure 3: The masks generated from different methods. From left to right are: SEEM, adaptive resolution SEEM (AdaSEEM), and AdaSEEM + Iter. 0 predictions. In (c), the new pseudo-label masks are highlighted with blue ellipses.
@@ 4.1 Experimental Settings
Dataset: We evaluate the proposed method on JHU-CROWD dataset [36], UCF-QNRF [17], ShanghaiTech [53] and UCF-CC-50 [16] datasets. The JHU-CROWD dataset is a comprehensive large-scale dataset, comprising 4,371 images. It is divided into three subsets: 2,722 images for training, 500 for validation, and 1,600 for testing. The UCF-QNRF dataset includes 1,535 images, with 1,201 designated for training and 334 for testing. The ShanghaiTech dataset is split into two parts: the ShanghaiTech A, which contains a total of 782 images, divided into 482 for training and 300 for testing, and the ShanghaiTech B, which includes 1,116 images, with 716 used for training and 400 for testing. UCF-CC-50 contains 50 grayscale images and we use 5-fold cross-validation in experiments.
@@
The parameters $\omega$ and $\beta$ in our loss function play crucial roles in optimizing performance. These parameters are set to 100 and 0.01, respectively, based on the ablation study in Figures 11 and 11. The threshold $\tau$ in AdaSEEM is set to 0.3 according to the experimental result shown in Figure 11.
@@ 4.2 Comparison with State-of-the-art Methods
To assess the effectiveness of our proposed method, we conducted a thorough evaluation by comparing it with both state-of-the-art unsupervised and supervised methods. The results of this comparison are detailed in Table 1. First, our proposed method outperforms other unsupervised methods in terms of MAE and MSE, and the margin of improvement is significant. This underscores the effectiveness of our approach in addressing the challenges inherent to unsupervised counting. Second, the comparison also reveals that the performance in Iter. 1 of our method is better than in Iter. 0 across all datasets. This improvement validates the effectiveness of our iterative pseudo-label generation strategy. By refining the pseudo-labels, the model is able to achieve more accurate and reliable counting results. We also compare the proposed method with cross-domain methods, which train on a source dataset and test on the target dataset, in Table 1 and achieve superior performance for most of the cases. Finally, the proposed method also compares favorably with some classic supervised methods. However, there is still considerable room for improvement, particularly in handling densely crowded datasets.
@@ 4.3 Ablation Study
Adaptive resolution SAM We conducted counting experiments to evaluate the effectiveness of AdaSEEM, and the results are presented in Table 2. The performance of SEEM on its own was found to be the least effective, indicating that directly using SEEM is not optimal due to the omission of many small individuals in high-density areas. However, with the implementation of the proposed adaptive resolution strategy, there was a noticeable performance improvement, especially for the high-density dataset UCF-QNRF. In Figure 4, we can also observe significant improvement in high-density images when using AdaSEEM. This improvement underscores the efficacy of the adaptive resolution strategy in accurately segmenting small persons in densely populated regions. The effectiveness of this approach is further confirmed in Figure 3(a, b), where more individuals are segmented with the AdaSEEM compared to the base model. These findings collectively highlight the significance of the adaptive resolution strategy in enhancing the segmentation capabilities of SEEM in complex crowd counting scenarios.
@@
Ours (Iter. 0) 125.4 226.7 195.9 343.0
AdaSEEM + Iter. 0 323.4 470.8 347.5 612.1
Ours (Iter. 1) 122.8 217.8 181.2 304.7
@@
Figure 9: Error v.s. Iter.
@@
Iterative pseudo-label generation To enhance mask quality, we introduce an iterative pseudo-labels generation approach, leveraging the trained counting network. First, we predict individual locations in training images, considering local maxima above a threshold as potential person localizations, following [40]. These predicted locations are then used as prompts for SEEM segmentation, effectively localizing new people in dense areas. As Figure 3 illustrates, this method detects more people, evidenced by the increased mask count. Performance comparisons in Table 2 show marked improvements with iterative pseudo-label generation by comparing “AdaSEEM” and “AdaSEEM + Iter. 0”. Moreover, “Ours (Iter. 1)”, a newly trained counting network with these refined masks outperforms the prior iteration, confirming the method’s efficacy.
To determine the optimal number of iterations, we experimented on ShanghaiTech A, with results depicted in Figure 11. The findings indicate that peak performance is attained at the second iteration, after which the performance converges. Consequently, we opted for two iterations in subsequent experiments.
Loss hyperparameters Figures 11 and 11 show the ablation studies for different values of the loss hyperparameters, $\omega$ and $\beta$. Figure 11 and 11 shows the ablation study for $\tau$ and $K$.
Localization performance We further evaluate the localization performance of the proposed method on UCF-QNRF. The performance of our proposed unsupervised method was benchmarked against existing supervised methods, filling a gap as there were no comparable unsupervised crowd localization methods. Despite the lack of manual labeling during training, our method demonstrated commendable precision, which outperforms several supervised counterparts, as shown in Table 3. The recall of our method is lower than supervised approaches but can be improved significantly using the 1st iteration training, which confirms that more missed people are detected and pseudo-labeled. While the overall localization performance of the proposed method is still limited and falls short of the state-of-the-art supervised methods, the results are promising, particularly considering the absence of manual labels.
@@
Ours (Iter. 0) None 0.777 0.101 0.456
Ours (Iter. 1) None 0.677 0.263 0.476
@@ 6 Conclusion
Acknowledgments
This work was supported by a Strategic Research Grant from City University of Hong Kong (Project No. 7005665).